OpenStudy (anonymous):

Evaluate each expression if r=-2, n=-3 and t=3. 1. |8-t|+3 2. |t-3| -7 3. 5+|3-n| 4. |r+n|-7 5. |n-t|+4 6. -|r+n+t|

OpenStudy (whpalmer4):

So this is just arithmetic, and knowledge that |x| = x if x > 0 or -x if x < 0. For example, the first one: \[\left| 8-t \right| + 3 = \left| 8-3 \right| + 3 = 5 + 3 = 8\]

OpenStudy (anonymous):

Okay thank you (: